OpenStudy (anonymous):

hello help me solve this equation by the linear method. (1-x ²) y '+ xy = x

OpenStudy (blurbendy):

Does it have to be the linear method?

OpenStudy (loser66):

first ODE?

OpenStudy (anonymous):

first do this ((1-x ²) dy / dx / (1-x ²)) + (x / (1-x ²)) = (x / (1-x ²)) clears dy / dx dy / dx + ((x, y) / (1-x ²)) = (x / (1-x ²)) x ² is removed?

OpenStudy (anonymous):

\[(1-x^2)y'+xy=x\\ y'+\frac{x}{1-x^2}y=\frac{x}{1-x^2}\] Integrating factor: \[\mu(x)=\exp\left(\int\frac{x}{1-x^2}~dx\right)\\ \mu(x)=\exp\left(\ln\left|\frac{1}{\sqrt{1-x^2}}\right|\right)\\ \mu(x)=\frac{1}{\sqrt{1-x^2}}\] \[\frac{1}{\sqrt{1-x^2}}y'+\frac{x}{\left(1-x^2\right)^{3/2}}y=\frac{x}{\left(1-x^2\right)^{3/2}}\\ \frac{d}{dx}\left[\frac{x}{(1-x^2)^{3/2}}y\right]=\frac{x}{(1-x^2)^{3/2}}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\vdots \] Note that the equation is also separable, as I'm sure @blurbendy was thinking. \[(1-x^2)y'+xy=x~\iff~\frac{1}{1-y}~dy=\frac{x}{1-x^2}~dx\]